# A Geometric Substructure for Quantum Dynamics

### v4

@@ 5 Interaction with the gravitational field
As mentioned in our opening remarks there has been extensive discussion over many years of attempts to quantise the theory of general relativity, and more recently of the possibility of “gravitising” quantum theory as an alternative approach to resolving the disconnect between the two theories, each of which boasts major successes in its own domain. The identification of a Riemannian space underlying quantum dynamics suggests a different resolution of this problem, one which treats quantum theory and the theory of relativity on a more equal footing, and we are led to make the following final conjectures:
$\bullet$ The geometric space $S$ underlying a given quantum system can be considered jointly with space-time carrying the local gravitational field, with combined coordinates $(z^{i},x^{\mu})$, for $i=1\,,2\,,\dots$ and $\mu=0\,,1\,,2\,,3\,,$ and combined metric tensor and infinitesimal distance-squared
@@
$\bullet$ A quantum measurement is labelled not only by the coordinates $z^{i}$ of a corresponding point in $S$, but also by the space-time coordinates $x^{\mu}$ of the point or points in space-time at which it occurs.
@@
in particular during measurements on the quantum system. This implies that not only does the local gravitational field interact with the measurement process, the changing entropy content of the space $S$ during measurements can alter the local gravitational field. In short, changing entropy at the quantum level can be an unexpected source of gravitational field strength – a kind of “dark energy.”
@@ 6 Concluding remarks
The simple substructure identified in Sec. 2 provides a new way of thinking about quantum dynamics and measurements without suggesting any new observable effects. On the other hand, the generalizations speculated upon in the following sections may have far-reaching and important implications for physics. Further study is encouraged. Acknowledgement: The author is grateful for the mentorship provided over many years by H.S. Green, A.O. Barut and L. Bass.

### v5

@@ 5 Interaction with the gravitational field
As mentioned in the Introduction there has been extensive discussion over many years of attempts to quantise the theory of general relativity, and more recently of the possibility of “gravitising” quantum theory as an alternative approach to resolving the disconnect between the two theories, each of which boasts major successes in its own domain. The identification of a Riemannian space underlying quantum dynamics suggests a different resolution of this problem, one which treats quantum theory and the theory of relativity on a more equal footing, and leads to the following final proposals:
[…]
@@
$\bullet$ Interactions of quantum systems with their environment are labelled not only by the coordinates $z^{i}$ corresponding points in $S$ but also by the space-time coordinates $x^{\mu}$ of the point or points in space-time at which such interactions occur.
@@
in particular during measurements and perhaps also during more general interactions between quantum systems and their environment. This implies that the local gravitational field interacts with measurement processes in particular and may influence quantum dynamics more generally – see for example [13] for related discussions in other contexts. It implies also that the changing entropy content of the space $S$ during such processes can alter the local gravitational field. In short, changing entropy at the quantum level can be an unexpected source of gravitational field strength – a kind of “dark energy.” How this new type of interaction affects Einstein’s equation for the gravitational field and equations governing change to the structure of $S$ must be the subject of further study.
@@ 6 Concluding remarks
The simple substructure identified in Sec. 2 provides a new way of thinking about quantum dynamics and measurements without suggesting any new observable effects. On the other hand, the generalizations suggested in the following sections would surely have far-reaching and important implications for physics. Further study is encouraged.
Acknowledgement: The author is grateful for the mentorship provided over many years by H.S. Green, A.O. Barut and L. Bass., and thanks W.M.J. Bracken, D. Ellinas and P.D. Jarvis for helpful comments.